ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57 Vaccaro, Michelangelo, Roma, Italy Unicity of the topological type of the simply connected 3manifolds. This talk presents ideas how to prove the Poincaré conjecture in dimension